OpenStudy (anonymous):

Given the exponential equation 4x = 64, what is the logarithmic form of the equation in base 10? x = log base 2 of 64, all over log base 2 of 4 x = log base 2 of 4, all over log base 2 of 64 x = log base 10 of 64, all over log base 10 of 4 x = log base 10 of 4, all over log base 10 of 64

OpenStudy (er.mohd.amir):

x=64/4 taking log base 2 log x=log(64/4) log x= log 64 -log 4 log x=6 log 2-2 log 2